OpenStudy (anonymous):

'When rounded to the nearest tenth, what are the points of intersection of the system y equals x squared minus three x minus three, y equals x plus one question mark Graph of the parabolic function y equals x squared minus three times x minus three and the linear function y equals x plus one. (−0.8, 0.2) and (4.8, 5.8) (−1, 1) and (5, 6) (0.2, −0.8) and (6, 5) (−0.9, 0.1) and (4.9, 5.9)'

jhonyy9 (jhonyy9):

@kaaare_ can you coming back please to cooperate please ?- so than i will try help you

jhonyy9 (jhonyy9):

y=x^2 -3x -3 y=x+1 - do you can solving this system please ?

jhonyy9 (jhonyy9):

by substitution is the best easy way - so substitute the value of y = x+1 in place of y inside first equation and will get x+1=x^2 -3x -3 and now solve it for x - ok ?

jhonyy9 (jhonyy9):

courage please this is easy - and i m here to help you - ok. ?

OpenStudy (anonymous):

so would it be (0.2 ,-0.8)?

jhonyy9 (jhonyy9):

i dont know - please write your calculi here

jhonyy9 (jhonyy9):

first solve this equation for x

OpenStudy (anonymous):

x=2(1+sqrt2) , x=2-2sqrt2

jhonyy9 (jhonyy9):

why you dont like write here your calculi for i can check them ?

OpenStudy (anonymous):

used calculator idk how to do this.

jhonyy9 (jhonyy9):

so first solve this for x x+1=x^2 -3x -3 add to both sides (-x-1) - so what will get ?

jhonyy9 (jhonyy9):

comen i will drive you step by step till the end of - ok. ?

jhonyy9 (jhonyy9):

so than you dont like cooperate - bye